Write an expression describing a transverse waves traveling along the string in the+ x direction wave length 11.4 cm,frequency 385hz ,and amplitude 2.13


1
Expert's answer
2020-12-25T14:06:05-0500

Explanations & Calculations


  • The relationship for the displacements of a particle on a transverse sinusoidal wave moving in +x direction is given by,

"\qquad\qquad
\begin{aligned}
\small y_{(x,t)}&= \small A\sin (kx-\omega t)\\

\end{aligned}"


  • A = amplitude of the wave
  • "\small k=\large \frac{2\pi}{\lambda}" and "\small \omega =2\pi f=2\pi \large\frac{v}{\lambda}"
  • "\small f=" frequency and "\small v=" speed of the wave


  • Substituting the given values,
  • "\small k=\large\frac{2\pi}{0.114m}=\small 55.12\,m^{-1}"
  • "\small \omega = 2\pi\times 385s^{-1}=2419.03\,s^{-1}"


  • These yeild,

"\qquad\qquad
\begin{aligned}
\small\bold{ y_{(x,t)}}&= \small \bold{2.13(55.12x-2419.03t)}
\end{aligned}"


  • Additionally, the speed of a transverse wave on a string is given by

"\qquad\qquad
\begin{aligned}
\small v&= \small \sqrt {\frac{T}{\rho}}
\end{aligned}" : "\small \rho" is the mass per unit length/linear density

  • And also waves on a string obays the wave equation,

"\qquad\qquad
\begin{aligned}
\small \frac{\partial^2y}{\partial x^2}&= \small \frac{1}{v^2}\frac{\partial^2y}{\partial t^2}
\end{aligned}"